A force of 20N pulls two carts tied together with a string. The carts with masses 3kg and 4kg, respectively, run on a smooth level surface. Find :


a. The acceleration of the 4kg cart


b. The acceleration of the 3kg cart


c. The tension on the string connecting the two carts.

1
Expert's answer
2021-11-09T10:58:40-0500

(a)-(b)


"a_1=a_2=a,""F=(m_1+m_2)a,""a=\dfrac{F}{(m_1+m_2)}=\dfrac{20\ N}{(3\ kg+4\ kg)}=2.86\ \dfrac{m}{s^2}."

(c)

"F-T=m_1a,""T=F-m_1a=20\ N-3\ kg\times2.86\ \dfrac{m}{s^2}=11.42\ N."